ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 205, 2015

Achar, P. N. and S. Riche Constructible sheaves on affine Grassmannians and geometry of the dual nilpotent cone

247

Albert, M. H., N. Ruškuc and V. Vatter Inflations of geometric grid classes of permutations

73

Avitabile, M. and S. Mattarei Laguerre polynomials of derivations  109

Bastin, F., J. A. Conejero, C. Esser and J. B. Seoane-Sepúlveda Algebrability and nowhere Gevrey differentiability

127

Caranti, A. A module-theoretic approach to abelian automorphism groups

235

Chorny, B. Homotopy theory of relative simplicial presheaves  471

Conejero, J. A. See F. Bastin

Cotón, C. M. n. Secondary LS category of measured laminations  211

Esser, C. See F. Bastin Farkas, H. M. and J. Y. Kaminski Construction of period matrices

by algebraic techniques

453

First, U. A. General bilinear forms  145

Giudici, M., C. H. Li, Ṡeress and A. Thomas Characterising vertex-star transitive and edge-star transitive graphs

35

Kaminski, J. Y. See H. M. Farkas Li, C. H. See M. Giudici

Matsumoto, K. Full groups of one-sided topological Markov shifts  1

Mattarei, S. See M. Avitabile Matz, J. Bounds for global coefficients in the fine geometric expansion

of Arthur’s trace formula for GL(n)

337

Pfitzner, H. L-embedded banach spaces and measure topology  421

Riche, S. See P. N. Achar Ruškuc, N. See M. H. Alpert

Sarkar, J. Submodules of the Hardy module over the polydisc  317

Seoane-Sepúlveda, J. B. See F. Bastin Seress, ˙ See M. Giudici Thomas, A. See M. Giudici

Thomsen, K. Exact circle maps and KMS states  397

Vatter, V. See M. H. Albert

Weigel, T. Graded Lie algebras of type FP  185
Zhao, D. Schur elements of degenerate cyclotomic Hecke algebras  485